# Linguistic Generalizations are not Rules: Impacts on Evaluation of LMs

### v1

@@ 4 Compositionality in LM Evaluation
As computer coding languages became more and more widespread, rule-based semantics and syntax took root in linguistics. A Principle of Compositionality combined the two traditions. It states that the meaning of a sentence is determined by the meanings of the words and the syntactic rules used to combine them Montague (1970); Partee (1984); Dowty (1979); Jackendoff (1992); Fodor and Lepore (2002). This is a bottom-up process: syntactic rules combine words, which have determinant meanings. Context was not supposed to influence the interpretation of words in a top-down manner. Instead, downstream inferences were invoked to address the obvious fact that interpretation does depend on context. As Fodor and Pylyshyn (1988) state, “a lexical item must make approximately the same semantic contribution to each expression in which it occurs”. Yet they, like Carnap and Frege, acknowledge: “It’s uncertain exactly how compositional natural languages actually are” Fodor and Pylyshyn (1988).
The standard argument in favor of compositionality is outlined in White et al. (2024):
  1. 1.
@@
  2. 2.
    Sentences are generated by a relatively small set of syntactic rules that combine words Chomsky (1957). $\Rightarrow$ Meaning is determined by meaningful words and the syntactic rules used to combine them.
While people more or less agree on the meanings of new sentences in context, this does not entail that meaning is determined by rule-based algorithms operating on familiar words (contra [1]). People also generally agree on the interpretations of pointing gestures and novel words, and yet in each of these cases, the shared interpretations must be gleaned from non-linguistic context (in the case of pointing gestures), or from a combination of linguistic and non-linguistic context.
Similarly, the meanings of familiar collocations, compounds and idioms are not determined by general rules. For instance, a compositional rule involving set-intersection may be appealing for ‘<color term> noun’ combinations in the domain of artificial block worlds (e.g., a green cube is something that is both a cube and green). However, violations of such rules abound e.g., green tea is more yellow than green and Cambridge blue is actually green. Even more common are instances that evoke richer meanings than predicted by any algebraic rule: e.g., a green light implies that forward motion or progress is permitted, and a green card provides a path toward citizenship in the US.
Critically, people tend to mostly agree on the interpretations of utterances in context-dependent ways. Consider ‘the Persian cat is on the rug.’ If the goal of the speaker is only to find the furball, there need be no commitment to the cat being a thoroughbred Persian breed. Likewise, the speaker need not be committed to the cat being wholly on, rather than adjacent to, the rug. Or, comprehenders may appreciate the statement as ironic, if the cat is hairless. And the rug need not have been created or used as a rug for people to share the same intended interpretation of the sentence.
The second premise in the standard argument is also problematic. Rules massively over-generalize, which is why they were never used for production. That is, rules predict all manner of odd locutions Pawley and Syder (1983); Sag et al. (2002): e.g., ‘Meeting you is pleasing to me’; ‘The tall winds hit the afraid boy’; ‘Explain him the problem.’ Humans are sensitive to the frequencies of various types of word combinations and judge formulations unnatural whenever there exists a more conventional way to express the intended message in context (e.g., Goldberg, 2019).
@@
Compositionality benchmarks combine elements from one or another evaluation paradigm already described. Kim and Linzen (2020)’s compositional generalization challenge (COGS) tested whether models could formalize into formal semantics, any sentence generated by a small set of syntactic rules. They anticipated generalizations from sentences like ‘Jane gave the cake to John’ to ‘Jane gave John the cake.’ The models were found to perform poorly. Speakers’ choice between these two constructions is highly sensitive to information structure, dialect Bresnan and Ford (2010), and the relative frequencies and similarities of verbs witnessing in each version Goldberg (2019); Leong and Linzen (2024); Ambridge et al. (2014). Insofar as natural language is not amenable to logical representations, failures of LMs to map to such representations may be consistent with human interpretation.
@@
The robotic agents struggled to interpret the rule-based command, though it was appropriate in the narrow domain tested. Notably, the rule does not extend to a broader swath of language. For instance, unbounded actions are not countable, so if ‘twice’ appears at all, it is likely followed by a comparative phrase (e.g., ‘work twice as hard’) and a very different meaning than performing an action two times. Other cases require knowledge of specific combinations: ‘to think twice,’ which means ’to hesitate’ and ‘going twice’ tends to evoke the context of an auction. Familiar phrases with meanings not fully captured by compositional rules are common: By one estimate, we learn tens of thousands of them Jackendoff (2002). Importantly, we largely agree on their interpretations, even though each means something more or different than predicted simply by the words and their syntactic combination. In this way, phrasal combinations regularly involve subregularities or item-specific interpretations not predicted by a general algebraic rule.
Another example comes from the seemingly innocuous algebraic rule stated below:
If X is more Y than Z, then Z is less Y than Z, irrespective of the specific meanings of X, Y, and Z. (Dasgupta et al., 2020, p.5)
The rule is intended to capture that ‘Anne is more cheerful than Bob’ should both contradict ‘Anne is less cheerful than Bob’, and entail ‘Bob is less cheerful than Anne.’ NLI models that failed to draw these inferences were considered lacking. Yet natural language rarely relies on free variables. The content of X, Y, and Z matter. No one would infer that because Annex is more cheerfuly than carefulz, that ‘Carefulz is less cheerfuly than Annex.’ Perhaps more importantly, if a speaker uttered ‘Anne is higher than Bob and Bob is higher than Anne,’ listeners would likely infer either that Bob climbed above Anne in the time it took to utter the first clause or that Bob has been smoking.
@@ 5 Constructions
We have argued against the idea that natural language is generated or interpreted by symbolic rules, but we agree that speakers are generally able to agree on the meanings of new sentences well enough for communication to be successful. This section briefly explains how constructions offer an alternative with respect to each of the differences cited in Table 1. Language is generated by flexibly combining constructions, which comprise a rich and complex ConstructionNet for each language. These include words, but are far broader than the traditional lexicon, encompassing schemata larger than individual words as well (Table 2).
(Partially-filled) Words, Common and Rare Schemata
We use the term ‘construction’ to refer to a learned association between a formal pattern and a range of related functions. This simple definition treats words, idioms, rare and common grammatical patterns as constructions, and recognizes that each case may include open ‘slots’ (Table 2). Formal attributes may include phonology, lexical content, grammatical categories, word order, discontinuous elements, and/or intonation.
@@
Constructions’ functions vary widely: Constructions may convey rich, specific contentful meaning in the case of words, collocations, idioms. A plethora of other constructions are productive but constrained in a wide variety of semi-specific ways; argument structure constructions convey ‘who did what to whom’; discourse structuring constructions indicate which subparts of a sentence are at-issue or backgrounded. Constructions exist to ask questions, express surprise or disapproval, for example. Any level of construction can be associated with specific registers, genres, and/or dialects.
Types of Construction and Examples
Words: pregame; running; nevertheless; ago
@@
Collocations, fixed idioms: high winds; jump the shark
@@
Partially-filled argument structure constructions: give <recipient> a call
Argument structure constructions: Intransitive; Caused-motion; Double Object; Resultative
Discourse-structuring constructions: get-passive; information questions; it-clefts; relative clauses
@@
Language users are sensitive to the frequencies of constructions. For instance, the passive construction is far more frequent in Turkish than English and young Turkish speakers learn the construction far earlier than English-speaking children Slobin (1986). Constructions are also influenced by similarity: Instances of a construction prime instances of the same or closely related construction (e.g., Du Bois, 2014; Pickering and Ferreira, 2008).
@@
As shown in Table 2, syntax, semantics and morphology are interrelated rather than assigned to distinct levels. This is useful because even productive hierarchical constructions often include particular words and semantic constraints. For example, an English construction that implies real or metaphorical motion allows a wide range of verbs but requires the particular noun ‘way’ ( ’He charmed his way into the meeting’.)
@@
The broad definition of constructions as pairings of form and function, including words and grammatical patterns, is another difference between constructions and rules. Constructions are far richer and more plentiful than the class of rules is commonly envisioned to be. Constructions also do far more work than rules since they capture frequency information and contextual constraints, while rules are presumed to be context-free and uninfluenced by frequency.
@@
The open ‘slots’ of constructions are constrained in a wide variety of ways. For instance, the English double-object construction can appear with a wide range of verbs, but prefers simple verbs to those that sound Latinate (e.g., ’She told them something’ vs. ’She proclaimed them something’). The English comparative suffix ‘-er’ (e.g., ‘calmer’, ‘quicker’) is available for most single-syllable adjectives that allow a gradient interpretation, but it is not used with past participles adjectives (? ‘benter’).
@@
Consider ‘X is the new Y’. It is productive and can be used to create new utterances e.g., ‘Semiconductor chips are the new oil.’ As is typical of productive constructions, the generalization co-exists with several familiar instances (e.g., ‘50 is the new 40’; ‘Orange is the new black’). The construction is not an algebraic rule. Its slots, indicated by X and Y, are not variables that range freely over fixed syntactic categories. Instead, ‘X’ must be construed (playfully) as currently functioning in the culture as ‘Y’ used to. Therefore not all combinations of slot fillers make sense: (e.g., ? ‘Orange is the new oil’). Adding a parallelism constraint between X and Y is insufficient since ‘103 is the new 101’ also makes little sense. Finally, instances of the construction are not amenable to translations into formal logic, which would presumably treat ‘Orange is the new black’ as equivalent to ‘Black is the old orange,’ which does not conventionally evoke the same meaning.
@@ 6 Implications Beyond Natural Language
The current observations help make sense of LM behavior outside the domain of pure language. Even in domains that are rule-like by design, certain types of non-compositional behavior exist, likely due to their interface with natural language. For instance, LMs have been found unreliable at drawing the following inference, which the authors dubbed the reversal curse: “if ‘A is B’ […] is true, then ‘B is A’ follows by the symmetry property of the identity relation" (Berglund et al., 2023, p. 2).
@@
Why are LMs prone to the reversal curse? Although the quote above is stated in natural language, it does not apply to natural language sentences, which are actually rarely reversible. For example, ‘A mental illness is the same as a physical illness’ means something very different than ‘A physical illness is the same as a mental illness’ (see also Tversky, 1977; Talmy, 1975). Even simple conjunctions are not generally reversible in natural language. For instance, ‘night & day’ and ‘day & night’ are both acceptable, but their interpretations differ: the former conveys a stark contrast (e.g., ‘as different as night and day’), the latter suggests a relentless activity or process (e.g., ‘he worried day and night’). In summary, it is perhaps reasonable to expect truly symmetric knowledge to be reversible. But LMs are trained on natural language and natural language utterances are not symmetric.
Human reasoning depends on the context in which performance is tested Klauer et al. (2000); Wason (1968); Tversky and Kahneman (1974) and how instructions are formulated Evans et al. (1994). Lampinen et al. (2024) find that LMs and humans are influenced by semantic context in similar ways (see also McCoy et al., 2023). Even math is not fully rule-compositional when equations are intended for communication: for instance, 2+2=42242+2=42 + 2 = 4 means something different than 4=2+24224=2+24 = 2 + 2 Mirin and Dawkins (2022), a preference picked up on by LMs Boguraev et al. (2024).
7 New Directions
@@
Natural languages involve complex and context-sensitive systems of constructions, which vary from being wholly fixed to highly abstract and productive. Constructions are combined when a unit, potentially itself composed of constructions, fills a slot in another construction. Viewing language as a system of constructions rather than words and rules may fundamentally change how the successes and failures of models are construed, and new goals and questions come into focus.
Balancing Constructions and Rules
Coding languages are compositional by design. They are unambiguous with variables filled by any instance of a clearly defined and general type. Accordingly, increasing the proportion of code in pretraining improves performance on tasks that rely on rule-based compositionality such as logic and math (e.g., Kim et al., 2024; Madaan et al., 2022). Yet Petty et al. (2024) report that adding code to pretraining data hurts performance on naturalistic language processing including tasks involving the English passive Mueller et al. (2022) and BigBench’s Implicatures and CommonMorpheme tasks Srivastava et al. (2023). Sprague et al. (2024) report that while Chain-of-Thought prompting had been believed helpful across-the-board, performance only improves on problems that require algorithmic reasoning. Thus, adding code to pre-training or using CoT prompts benefit tasks that are designed to be rule-compositional but may be detrimental to to natural language tasks.
Better Datasets
@@
Rather than using abstract rules to generate stimuli for natural language benchmarks, ecologically valid stimuli may be more usefully collected or adapted from natural corpora and then normed for naturalness and plausibility. Since human judgments are highly context-dependent, benchmark tasks should ideally also vary contexts systematically (see, e.g., Ross et al., 2024).
It is also important to avoid inadvertently training human coders to give the type of responses only suitable in logic or coding classes. If people are instructed to interpret ‘red X’ as ’X that is red for any X,’ they can do so. Yet in natural contexts, people understand that red grapefruits are closer to pink, red hair is more orange, a red book may be about communism, and crossing a red line may have consequences. Finally, a variety of items, participants, and contexts ought to be valued as much as a variety of models.
Probing for Constructions
Recognizing that LMs are already extremely skilled at producing and responding to several natural languages allows for a shift in research agendas. The interesting question may now be not if, but how LMs achieve such remarkable skills. We can now also ask how well LMs capture appropriately nuanced interpretations and relationships among constructions. New ways of probing LMs make this possible and this toolkit will only grow.
As is familiar from the lexicon, constructions are related to one another because human memory is highly associative. Misra and Mahowald (2024) have demonstrated that even when all instances of a rare non-compositional construction are ablated from training data, non-trivial learning of the construction remains, enabled by the presence of related constructions in training. Another type of relationship among constructions are the relationships between conventional instances and productive generalizations. Nearly every productive construction co-exists with at least a few formulaic instances, and LMs offer ways of testing relationships among instances that give rise to productive generalizations. Other recent work includes Weissweiler et al. (2022), who found LMs reliably discriminate instances of the English Comparative Correlative from superficially similar expressions. Tayyar Madabushi et al. (2020) tested a dataset of automatically induced constructions and reported that BERT Devlin et al. (2019) could determine whether two sentences contained instances of the same construction. However, Zhou et al. (2024) found LMs failed to distinguish entailment differences between the causal excess construction (e.g., ‘so heavy that it fell’) and two structurally similar constructions (‘so happy that she won’; ‘so certain that it rained’). Similarly addressing challenging construction semantics, Weissweiler et al. (2024) showed LMs to struggle with the meaning of the caused-motion construction.
Tseng et al. (2022) showed that LMs gradiently predict appropriate slot fillers. Li et al. (2022) probed RoBERTa’s implicit semantic representations of four argument structure constructions and found similarities in behavior in the model and a sorting task done by humans. Potts (2023) found that despite its rarity, LMs acquire the Preposing in PP construction Huddleston and Pullum (2002).

### v2

@@ 4 Compositionality
As computer coding languages became more and more widespread, rule-based syntax and semantics took root in linguistics. A Principle of Compositionality states that the semantics of a sentence is determined by the meanings of the words and the syntactic rules used to combine them Montague (1970); Partee (1984); Dowty (1979); Jackendoff (1992); Fodor and Lepore (2002). It is intended to be a bottom-up process: syntactic rules combine words, which have determinant meanings. Fodor and Pylyshyn (1988) make this clear: “a lexical item must make approximately the same semantic contribution to each expression in which it occurs”. That is, context may not influence the interpretation of words in a top-down manner; therefore downstream inferences are required to address the fact that interpretations do depend on context. Realizing this, like Carnap and Frege before him, Fodor and Pylyshyn (1988) acknowledge: “It’s uncertain exactly how compositional natural languages actually are.”
Nonetheless, compositionality is often taken as a truism, based on the standard argument for it summarized below.
@@
One can agree with the premise without accepting its consequent.
In particular, people generally agree on the interpretations of pointing gestures and novel words as well as sentences, and yet shared interpretations must be gleaned from non-linguistic context in the case of pointing gestures, and from a combination of linguistic and non-linguistic context in the case of novel words. Shared interpretation of sentences likewise comes in part from linguistic and non-linguistic context.
Consider the sentence, ‘the Persian cat is on the mat.’ If the speaker’s goal is simply to help someone find the furball, there need be no commitment to the cat being a thoroughbred Persian breed nor to the cat being wholly on, rather than adjacent to, the mat. Or, comprehenders may appreciate the statement is ironic if the cat is hairless.
Cases that might seem amenable to a rule often turn out to require a good deal of item-specific memory. For instance, a compositional rule involving set-intersection may seem appealing for ‘<color term> noun’ combinations in the domain of artificial block worlds (e.g., a green cube is something that is both a cube and green). However, violations of such rules abound: green tea is more yellow than green, and Cambridge blue is actually green. Even more common are instances that evoke richer meanings than predicted by any algebraic rule: e.g., a green light implies that forward motion or progress is permitted, and a green card provides a path toward citizenship in the US (or ought to). The meanings of familiar collocations and idioms are typically not fully determined by general rule.
Another issue is that rules massively over-genenerate, which is why they are not used for production. That is, rules predict all manner of odd locutions Pawley and Syder (1983); Sag et al. (2002): e.g., ‘Meeting you is pleasing to me’; ‘The tall winds hit the afraid boy’; ‘Explain them the problem.’ Humans are sensitive to the frequencies of various types of word combinations and judge formulations unnatural whenever there exists a more conventional way to express the intended message in context (e.g., Goldberg, 2019).
@@
Compositionality benchmarks combine elements from syntactic and semantic evaluations. Kim and Linzen (2020)’s compositional generalization challenge (COGS) tested whether models could translate any sentence generated by a small set of syntactic rules into formal semantics. For instance, trained on representations of ‘the girl,’ ‘the cat,’ ‘the hedgehog,’ ‘the cat loves the girl,’ and ‘the hedgehog sees the cat,’ and so on, the model was tested on how well it predicted a formal semantic representation for ‘The girl loves the hedgehog.’ However, note that if ‘mosquitoes’ is substituted for ‘the cat,’ different interpretations of ‘love’ are evoked (‘Mosquitoes love the girl’ vs. ‘The girl loves mosquitoes’), not to mention markedly different degrees of plausibility. The authors also anticipated generalizations from sentences like ‘Jane gave the cake to John’ to ‘Jane gave John the cake,’ and the models were found to perform poorly. Yet the two sentences differ in terms of information structure Bresnan and Ford (2010), and the relative frequencies and similarities of verbs witnessed in each version Leong and Linzen (2024); Ambridge et al. (2014); Hawkins et al. (2020). Insofar as natural language is not amenable to logical representations, failures of LMs to map to such representations may be consistent with human interpretation.
@@
The robotic agents struggled to interpret the rule-based command, though it was appropriate in the narrow domain tested. Notably, the rule does not extend to a broader swath of language. For instance, unbounded actions are not countable, so if ‘twice’ appears at all, it is likely followed by a comparative phrase (e.g., ‘work twice as hard’) and a very different meaning than performing an action two times. Other cases require knowledge of specific combinations: ‘to think twice’ means ’to hesitate’ and ‘going twice’ tends to evoke the context of an auction. Familiar phrases with meanings not fully captured by compositional rules are common: By one estimate, we learn tens of thousands of them Jackendoff (2002). Importantly, we largely agree on their interpretations, even though each means something more or different than predicted simply by the words and their syntactic combination. In this way, phrasal combinations regularly involve subregularities or item-specific interpretations not predicted by a general algebraic rule.
Another example comes from the seemingly innocuous algebraic rule: “If X is more Y than Z, then Z is less Y than Z, irrespective of the specific meanings of X, Y, and Z” (Dasgupta et al., 2020, p.5). This is meant to capture that ‘Anne is more cheerful than Bob’ should both contradict ‘Anne is less cheerful than Bob’, and entail ‘Bob is less cheerful than Anne.’ NLI models that failed to draw these inferences were considered lacking. Yet natural language rarely relies on free variables. The content of X, Y, and Z matters. No one would infer that because Annex is more cheerfuly than carefulz, that ‘Carefulz is less cheerfuly than Annex.’ Perhaps more importantly, if a speaker uttered ‘Anne is higher than Bob and Bob is higher than Anne,’ listeners would likely infer either that Bob climbed above Anne in the time it took to utter the first clause or that Bob has been smoking. We have so far argued against symbolic abstract rules for natural languages. Below we suggest a positive alternative.
@@ 5 Constructions
This section briefly explains an alternative to positing a set of abstract symbolic rules, namely a network of interrelated constructions. Constructions differ from rules in each dimension cited in Table 1 and briefly explained below. For more comprehensive introductions, see Hoffmann and Trousdale (2013) and Hoffmann (2022).
(Partially-filled) Words, Common & Rare Schemata
A ‘construction’ is any learned association between a formal pattern and a range of related functions. This simple definition treats words, idioms, rare and common grammatical patterns as constructions. Formal attributes may include phonology, lexical content, semantic or formal categories, word order, discontinuous elements, and/or intonation. Any construction may include one or more constrained open ‘slots’ (Table 2).
@@
Constructions’ functions vary widely: Constructions convey rich, specific, contentful meaning in the case of words, collocations, and idioms. A plethora of other constructions are productive but constrained in a wide variety of semi-specific ways; argument structure constructions convey ‘who did what to whom’; discourse structuring constructions indicate which parts of an utterance are at-issue or backgrounded. A range of constructions exist to ask questions, express surprise or disapproval, greetings or gossip. Any level of construction can be associated with specific registers, genres, and/or dialects.
Construction Type Examples
Words pregame; running; nevertheless; ago
@@
Collocations, fixed idioms high winds; jump the shark
@@
Partially-filled argument structure constructions give <recipient> a call
Argument structure constructions Intransitive; Caused-motion; Double Object; Resultative
Discourse-structuring constructions get-passive; information questions; it-clefts; relative clauses
@@
Language users are sensitive to the frequencies of constructions. For instance, the passive construction is used far more frequently in Turkish than English and young Turkish speakers learn the construction far earlier than English-speaking children Slobin (1986). Constructions are also influenced by similarity: Instances of a construction prime instances of the same or closely related construction (e.g., Du Bois, 2014; Pickering and Ferreira, 2008).
@@
As shown in Table 2, syntax, semantics, and morphology are interrelated rather than assigned distinct levels. This is useful because even productive hierarchical constructions often include particular words and semantic constraints. For example, an English construction that implies real or metaphorical motion allows a wide range of verbs but requires the particular noun ‘way’ (‘He charmed his way into the meeting.’).
@@
The broad definition of constructions as pairings of form and function, including words and grammatical patterns, is another difference between constructions and rules. Constructions are far richer and more plentiful than the class of rules is commonly envisioned to be, as they combine aspects of form and function. Constructions also do far more work than rules since they capture frequency information and contextual constraints, while rules are presumed to be context-free and uninfluenced by frequency (e.g., Pinker, 1999).
@@
The open ‘slots’ of constructions are constrained in a wide variety of ways. For instance, the English double-object construction can appear with a wide range of verbs, but prefers simple verbs to those that sound Latinate (e.g., ‘She told them something’ vs. ¿She proclaimed them something’). The English comparative suffix ‘-er’ (e.g., ‘calmer’, ‘quicker’) is available for most single-syllable adjectives that allow a gradient interpretation, but it is not used with past participle adjectives (? ‘benter’).
@@
Consider ‘X is the new Y’. It is productive and can be used to create new utterances, e.g., ‘Semiconductor chips are the new oil.’ As is typical of productive constructions, the generalization co-exists with several familiar instances (e.g., ‘50 is the new 40’; ‘Orange is the new black’). The construction is not an algebraic rule: Its slots, indicated by X and Y, are not variables that range freely over fixed syntactic categories. Instead, ‘X’ must be construed (playfully) as currently functioning in the culture as ‘Y’ used to. Therefore, not all combinations of slot fillers make sense: (e.g., ? ‘Orange is the new oil’). Adding a parallelism constraint between X and Y is insufficient since ‘103 is the new 101’ would also require an unusual context to make sense. Finally, instances of the construction are not amenable to a general compositional rule, nor can it be translated into formal logic, which would presumably treat ‘Orange is the new black’ as equivalent to ‘Black is the old orange,’ which does not conventionally evoke the same meaning.
@@ 6 Implications Beyond Natural Language
Outside of natural language, even in domains that are rule-like by design, certain types of non-rule-based interpretations exist, likely due to the interface from natural language. For instance, LMs have been found unreliable at drawing the following inference, which the authors dubbed the reversal curse: “if ‘A is B’ […] is true, then ‘B is A’ follows by the symmetry property of the identity relation" (Berglund et al., 2023, p. 2).
@@
Why are LMs prone to the reversal curse? Although the quote above is stated in natural language, it does not apply to natural language sentences, which are rarely reversible with the same interpretation: e.g., ‘A mental illness is the same as a physical illness’ means something very different than ‘A physical illness is the same as a mental illness’ (see also Tversky, 1977; Talmy, 1975). Even simple conjunctions are not generally reversible in natural language. For instance, ‘night & day’ and ‘day & night’ are both acceptable, but their interpretations differ: the former conveys a stark contrast (e.g., ‘as different as night and day’), the latter suggests a relentless activity or process (e.g., ‘he worried day and night’). In summary, it is perhaps reasonable to expect truly symmetric knowledge to be reversible. But LMs are trained on natural language, which is not symmetric.
7 New Directions for Evaluation
@@
Natural languages involve complex and context-sensitive systems of constructions, which vary from being wholly fixed to highly abstract and productive (recall Table 2). Constructions are combined when a unit, potentially itself composed of constructions, fills a slot in another construction. Viewing language as a system of constructions rather than words and rules may fundamentally change how the successes and failures of models are construed, and new goals and questions come into focus. The perspective on language motivates a move away from certain traditional evaluation practices. Specifically, we recommend the following.
@@
When possible, use a variety of naturalistic sentences rather than sentences generated by a template that presupposes grammatical rules with interchangeable vocabulary items. The idea that sentences can be constructed by subbing random lexical items into templates often misses lexical subtleties that are an important part of natural language. Instead, ecologically valid stimuli can be collected or adapted from natural corpora and normed for naturalness and plausibility. Since human judgments are highly context-dependent, benchmark tasks should also vary contexts systematically (see, e.g., Ross et al., 2024; Parrish et al., 2021).
Collect human assessments that allow for gradient context-sensitive interpretations that appeal to learned constructions. Evaluating LM competence on individual constructions requires assessing both acceptability judgments and interpretations from humans to make fair comparisons.
We also need to be sensitive to the implications people and LMs draw from instructions and the testing context. Do not give instructions to human-evaluators (or models) in ways that make the results a foregone conclusion. If people are instructed to interpret ‘red X’ as ‘X that is red for any X,’ they are capable of doing so; this may reflect the instructions, not their natural intuitions. In natural contexts, people understand that red grapefruits are closer to pink, red hair is more orange, a red book may be about communism, and crossing a red line may have consequences.
The items included in testing also influence interpretations by people and models, by providing context. For instance, if certain pairs of items are mismatched (e.g., “The cup is green.” “The cup is blue.”) while others are matched, people can infer which ones are intended to be contradictory. LMs, like people, are now capable of generalizing by rule when tasked to do so. For instance, Lampinen et al. (2025) found Gemini 1.5 Flash Team et al. (2024) avoided the reversal curse, achieving 100% accuracy, when the Berglund et al. (2023) dataset was provided to the LM as context.
A variety of items, participants, and contexts ought to be valued as much as a variety of models. Because linguistic meaning is deeply tied to local context, even seemingly similar sentences can have very different interpretations in ways that depend on context. Therefore, experiments that vary context are desirable.
The most interesting questions may no longer be whether LMs are skilled at producing and responding to natural languages, but how they achieve such remarkable skills. As is familiar from the lexicon, constructions comprise an interrelated network. We can now study relationships among constructions mechanistically. For instance, Misra and Mahowald (2024) have demonstrated that even when all instances of a rare non-compositional construction are ablated from training data, non-trivial learning of the construction remains, enabled by the presence of related constructions in training. Moreover, nearly every productive construction coexists with at least a few formulaic instances, and LMs offer ways to test various theoretical perspectives on the nature of those relationships. These and other newer ways of probing LMs are possible, and our toolkit will only grow.
We can investigate whether LMs distinguish subtle meaningful differences between constructions Tseng et al. (2022); Li et al. (2022); Weissweiler et al. (2022); Tayyar Madabushi et al. (2020); Devlin et al. (2019); Zhou et al. (2024). How they do so or when and why they fall short is another important direction.